A 8 kg block is at rest on a horizontal floor. If you push horizontally on the 8 kg block with a force of 20 N, it just starts to move.

(a) What is the coefficient of static friction?

(b) A 10.0 kg block is stacked on top of the 8 kg block. What is the magnitude F of the force, acting horizontally on the 8 kg block as before, that is required to make the two blocks start to move?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

ans a)if the block just starts moving then static friction force is equal to the applied force

  friction force=F (applied force)

    u.N =20    (here N is normal reaction =mg=80N)

     u=20/80=.25

ans b) F(min)=static friction force

                   =u.N             (N=mg =180N)

                   =0.25*180=45N